# algorithmic

### v2

@@ 1.5 Outline
In the next section, we begin with an introduction to the basics by revisiting the most important concepts from the standard RL setting, which are also naturally important in RLHF (Section 2). We then dive into the RLHF topic by outlining the most studied scenario of reward model learning from pairwise preferences. Using this introductory and illustrative example scenario, we explain the basic framework of RLHF alongside its three main components of (human) feedback, label collection (feedback acquisition), and reward model learning. These three main components will essentially form the structure of our survey. In Section 3, we turn our attention to the human feedback component and provide an overview of the different types of feedback as well as their key attributes. The important concepts in terms of label collection are then explained in Section 4, followed by learning the reward model in Section 5. We also discuss policy learning in Section 6, since some RLHF methods adapt standard RL training methods to the RLHF setting. Section 7 is devoted to an overview of recent progress on the theoretical side of RLHF, including approaches involving a theoretical guarantee, as well as theoretical insights into the relationship between standard RL and RLHF. Finally, Section 8 highlights some interesting practical applications of RLHF and the existing benchmarks before Section 9 concludes the survey by pointing out some possible avenues for future work.
@@ 2 Preliminaries
In this section, we recall the basic setting and the most important concepts of RL and RLHF. In the course of this review, we will fix the notation that will be used throughout the survey. We first introduce what is probably the most studied RLHF scenario, i.e., learning a reward model from binary trajectory comparisons. Based on this introductory and illustrative example scenario, we explain the basic framework of RLHF with its main components and briefly discuss the respective roles of these components in the learning process. We will also briefly touch on active learning, which strongly connects to the feedback collection component.
Notations
@@ 2.1 Reinforcement Learning
The transition function ${P}$ defines the dynamics of the environment: For any state $s$ and action $a$, the value ${P}(s,a)(s^{\prime})$, also sometimes denoted ${P}(s^{\prime}\mid s,a)$, is the probability of reaching the state $s^{\prime}$ after executing $a$ in $s$. In light of this, we will also refer to the transition function sometimes simply as the transition dynamics. For a given state and action, the transition is conditionally independent of all previous states and actions, which is known as the Markov property and the reason for the naming as an MDP. The value ${R}(s,a)\in\mathbb{R}$ provides an immediate evaluation after performing action $a$ in state $s$, which is also called the (instantaneous) reward. It is also quite possible that the instantaneous reward is 00 for some states, and one only receives a reward in specific states, for example, in so-called terminal states for which the transition function is zero. When both the state space ${\mathcal{S}}$ and the action space ${\mathcal{A}}$ are finite, we call the MDP a tabular MDP.
In an MDP, an $H$-step trajectory $\tau$ is a sequence of $\0H\in\mathbb{N}\backslash\{0\}$ pairs of state-action ending in a terminal state. Formally, it is given by $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H})$. Given $t_{0}\geq 0$ and $H^{\prime}\leq H$, we can define a segment $\sigma=(s_{t_{0}},a_{t_{0}},s_{t_{0}+1},a_{t_{0}+1},\dots,s_{H^{\prime}}$), which refers to a continuous sequence of steps within a larger trajectory. A trajectory $\tau$’s return ${R}(\tau)$ is the accumulated (discounted) rewards collected along this trajectory:
@@
Note that we here use the same notation for the return and the reward function, but both have different signatures (trajectory vs. state-action pair). We can also define the return $R(\sigma)$ of a segment $\sigma$ in a similar manner. The return is well defined even if the horizon $H$ is infinite as long as $\gamma<1$. If the MDP is a tabular MDP and any trajectory has finite length, i.e., $H$ is necessarily finite, we call the MDP finite, and otherwise infinite.
[…]
@@
(b) RLHF with reward modeling
Figure 1: Contrasting the standard RL setting with RLHF in its most common formulation, using a reward model. In each step, the policy commits to an action $a_{t}$ and receives the next state $s_{t+1}$ and either the true reward $r_{t+1}$ or an estimate $\hat{r}_{t+1}$ in return (symbolized by $\tilde{r}_{t+1}$). In contrast to the standard RL setting, the true reward function is not known in the RLHF setting but instead learned form human feedback. This reward learning process is decoupled from policy learning and can happen fully asynchronously. The dataset consists of a set of queries $q_{i}$ (e.g., pairs of trajectory fragments) and their labels $l_{i}$ (e.g., a preference for one of the fragments).
@@
The basic interaction loop is depicted in Fig. 1(a): The agent chooses an action $a_{t}\sim\pi(s_{t})$ based on its policy and the current state. As a consequence, the environment transitions into the new state $s_{t+1}\sim P(s_{t},a_{t})$, governed by the transition dynamics. The agent observes this new state and the reward $r_{t+1}\sim R(s,a)$, after which this interaction cycle is restarted.
@@
where the expectation is with respect to policy $\pi$, transition function ${P}$, and initial distribution $d_{0}$. To solve this problem, two families of RL approaches have been considered: model-based RL and model-free RL. The methods in the first family learn a model (i.e., ${P},{R}$) of the underlying MDP to help solve the RL problem, while the methods in the second directly try to obtain a good policy without learning an MDP model. The second family can be further decomposed into two main categories: value-based methods and policy search methods. In deep RL, both value functions and policies are approximated with neural networks.
@@
where $s_{0}=s$, and $a_{0}=a$ and in the expectation, $a_{h}\sim\pi(\cdot\mid s_{h})$ as well as $s_{h}\sim{P}(\cdot\mid s_{h-1},a_{h-1})$ for $h\in[H-1]$. A policy can be naturally designed from a $Q$-function by choosing an action in a greedy manner in each state: $\pi(s)=\arg\max_{a}Q(s,a)$. Note that for a deterministic optimal policy $\pi^{*}$ it holds that $J(\pi^{*})=\mathbb{E}_{d_{0}}[Q^{*}(s,\pi^{*}(s))]$.
@@ 2.2 Preference-Based MDPs
In contrast to standard RL as described in the previous section, RLHF does not assume that a reward signal is available. It instead assumes the existence of an oracle (e.g., human labeler) that can provide information about the reward in a specific indirect manner. More precisely, in the RLHF, the agent can make queries $q_{i}$ to the oracle, which in practice means asking for human feedback, and in response, the agent receives a label $l_{i}$, which in general gives a hint about the reward. In principle, the query can be made asynchronously to the actual conventional RL cycle. See Fig. 1(b) for an illustration.
@@ 2.3 Reward Learning
RLHF approaches can be divided into two categories, depending on whether a utility-based approach is used for reward modeling or an alternative criterion that is detached from a utility concept is used (Gilbert et al., 2016; Gilbert & Weng, 2016; Wirth et al., 2017). Most works fall into the first category, on which this overview focuses. Such approaches assume a human-dependent utility function that can be used as a reward function in order to apply standard RL methods. Next, we will describe the commonly used approach for reward learning for the common setting of binary trajectory comparisons.
@@
In the context of RL, since ${R}_{\psi}(\tau)=\sum_{h=0}^{H-1}\gamma^{h}R_{\psi}(s_{h},a_{h})$, (2) can then directly be used to train a function approximator (e.g., single or ensemble of neural network) to approximate $R$.
This entire modeling approach accommodates the case of a noisy or unreliable oracle, in which case the Bradley-Terry model can be understood as the generative model of the answers from the oracle (or labels provided by the human labeler). When the oracle is reliable, more direct methods based on preference elicitation to recover the reward function have been studied (Regan & Boutilier, 2009; 2011; Weng & Zanuttini, 2013; Gilbert et al., 2015; Sadigh et al., 2017; Wilde et al., 2018). In this survey, we will focus on the general case where the oracle may be noisy.
Note that in contrast to the typical way of preference learning, the learned reward function is used to train an RL agent and not directly to compare trajectories. This discrepancy in the objective function in the reward learning part and how the learned rewards are used may lead to suboptimal policies (Lindner et al., 2021b).
@@ 2.4 Reinforcement Learning from Human Feedback
In the RLHF setting as illustrated in Fig. 1(b), the learning agent needs to solve an RL task without having access to a reward function. To this end, the agent usually simultaneously learns an approximation of the reward function (via the assumed utility function) and an RL policy. Therefore, a generic RLHF algorithm consists of repeating two phases: (1) reward learning and (2) RL training. The first phase can itself be decomposed into two main steps: (i) generate queries to ask the oracle, (ii) train a reward function approximator with the answers provided by the oracle. The RL training part is more conventional and is usually directly based on running a deep RL algorithm using the currently trained reward function approximator.
Algorithm 1 Generic RLHF Algorithm in an Actor-Critic Scheme
@@
7:     Update $\psi$ using $\mathcal{D}$ (e.g., to maximize Eq. 2)
@@
This basic generic algorithm is summarized in Algorithm 1, where an off-policy actor-critic scheme is assumed to be used for the RL training part, but other RL policy learning approaches can, of course, also be used here. For an on-policy algorithm, such as PPO (Schulman et al., 2017), only the recently generated transitions are used for training. For a DQN-like algorithm, lines 9 to 11 would be replaced by a loop where transitions are generated by a behavior policy based on the current estimate of the $Q$-function (e.g., $\varepsilon$-greedy algorithm) and the $Q$ network is updated using mini-batches of transitions sampled from the replay buffer $\mathcal{D}$.
@@
  • The oracle may provide various types of feedback (see Section 3). The questions of what information some given feedback provides and how observed feedback can be exploited need to be answered (see Section 5).
  • Informative queries need to be generated to minimize the efforts of the oracle, which is crucial when it is a human (see Section 4). Active learning techniques (see next subsection) can be adapted to face this challenge.
  • The RL agent is actually trained in a non-stationary environment since the reward approximator is concurrently updated. The RL training part needs, therefore, to account for this factor, e.g., using non-vanishing learning rates (see Section 6).
  • There is also the question of how the agent’s performance can be meaningfully evaluated, especially if the reward function is not known (see Section 8).
  • Collecting feedback directly from humans introduces its own challenges, such as the question of a suitable user interface and the associated issues of delay between query and feedback observation, or the feedback variability and reliability (see Section 4). This may explain why many studies evaluate novel RLHF algorithms with simulated feedback.
A standard RL algorithm can be run in the RL training part, as done in most previous work in RLHF (although this may not be the best approach). This suggests that any improvements in a standard deep RL method (e.g., auxiliary losses (Gelada et al., 2019), planning in learned model (Hafner et al., 2020), curriculum learning (Narvekar et al., 2020), or data augmentation (Laskin et al., 2020; Lee et al., 2020; Lin et al., 2020b)) may potentially be transferred to the RLHF setting. In addition, most previous work in RLHF directly uses trajectories stored in replay buffer $\mathcal{D}$ to synthesize queries. An interesting research direction to explore in RLHF would be to specifically generate trajectories in order to be able to synthesize more informative queries (instead of only generating trajectories that are beneficial for RL training). This would lead to tackle a novel exploration-exploitation dilemma: Shall we visit state-action pairs that may be bad but may help better learn the reward function, or shall we visit state-action pairs that we currently think are good? This is further discussed in Section 5.5.3.
@@ 2.5 Active Learning
In active learning (Settles, 2012), the task is to strategically select data points for labeling to minimize the amount of labeled data required to achieve a desired level of performance, particularly valuable in scenarios like RLHF where labeling is costly. Unlike batch learning, where labeled data is predetermined, active learning empowers the learner to actively select the most informative unlabeled instances for labeling, maximizing the learning process with limited labeled data. We will only briefly introduce the active learning task here and then discuss the strategies for creating informative queries considered thus far in Section 4.
@@
At a high level, the main idea in active learning is to query data points to quickly reduce the epistemic (i.e., reducible) uncertainty about the predictions of the learned model, although other aspects can be important, such as the representativeness of the queried data points (see Wang et al. (2023a) for a survey). Two main representations are considered to describe this epistemic uncertainty: either using an ensemble of models or using a Bayesian representation. In both cases, a first basic approach selects queries using uncertainty-based criteria in order to focus on instances with high prediction uncertainty as measured by, e.g., variance or entropy computed over predictions. In contrast to the first approach, where the selection criteria are instance-based, a second approach considers criteria that may depend on all instances. Possible options are, for instance, expected model change, expected error reduction, or density-weighted uncertainty-based criteria. Here, the weights in the expectation or density allow us to take into account the distributional information about the instances and, therefore, to focus on the higher-density regions.
@@ 3 Feedback
Figure 2: Highlighting the components of the RLHF loop discussed in this section.
Feedback mechanisms are fundamental to the success of any RL system. In the standard setting as described in Section 2.1, the RL agents expect feedback in the form of scalar immediate rewards. These rewards are most commonly determined by a hand-engineered reward function, which can be used to evaluate any state-action combination. As discussed in Section 1.1, it is desirable to allow humans to refine behavior interactively through feedback instead of requiring them to pre-specify a reward function.
While a human could, in principle, assign rewards to each of the agent’s actions directly, thereby taking the role of the reward function, this is usually impractical for multiple reasons. The main challenge is the human effort required to provide rewards on a sufficiently regular basis, i.e., at least once per episode. In addition to that, directly integrating human rewards into the RL loop would require these rewards immediately, which impedes the learning pace while waiting for human feedback. Finally, the standard RL setting expects numeric state-action rewards, which is challenging to provide in a consistent manner.
In contrast to directly rewarding each of the agent’s actions, RLHF as discussed in this survey (see Section 1.3) harnesses indirect and asynchronous feedback methods. Such methods avoid the challenges of immediate numeric rewards and are also better aligned with human interaction patterns, resulting in improved learning progress and human user experience.
@@
The remainder of this section will start by discussing relevant attributes of feedback types that can be used to classify them (Section 3.1). We will then discuss common classes and examples of interactive feedback types (Section 3.2) as well as some non-interactive types that can serve as initializations (Section 3.3).

### v3

@@ 1.4 Outline
In the next section, we begin with an introduction to the basics by revisiting the most important concepts from the standard RL setting, which are also naturally important in RLHF (Section˜2). We then dive into the RLHF topic by outlining the most studied scenario of reward model learning from pairwise preferences. Using this introductory and illustrative example scenario, we explain the basic framework of RLHF alongside its three main components of (human) feedback, label collection (feedback acquisition), and reward model learning. These three main components will essentially form the structure of our survey. In Section˜3, we turn our attention to the human feedback component and provide an overview of the different types of feedback as well as their key attributes. The important concepts in terms of label collection are then explained in Section˜4, followed by learning the reward model in Section˜5. We also discuss policy learning in Section˜6, since some RLHF methods adapt standard RL training methods to the RLHF setting. Section˜7 is devoted to an overview of recent progress on the theoretical side of RLHF, including approaches involving a theoretical guarantee, as well as theoretical insights into the relationship between standard RL and RLHF. Finally, Section˜8 highlights some interesting practical applications of RLHF and the existing benchmarks before Section˜9 concludes the survey by pointing out some possible avenues for future work.
@@ 2 Preliminaries
In this section, we review the basic setting and the most important concepts of RL and RLHF. In the course of this review, we will establish the notation that will be used throughout the survey. We first introduce what is probably the most studied RLHF scenario, i.e., learning a reward model from binary trajectory comparisons. Based on this introductory and illustrative example scenario, we explain the basic framework of RLHF with its main components and briefly discuss the respective roles of these components in the learning process. We will also briefly touch on active learning, which strongly connects to the feedback collection component.
Notation
@@ 2.1 Reinforcement Learning
The transition function P{P} defines the dynamics of the environment: For any state $s$ and action $a$, the value ${P}(s,a)(s^{\prime})$, also sometimes denoted ${P}(s^{\prime}\mid s,a)$, is the probability of reaching the state $s^{\prime}$ after executing $a$ in $s$. In light of this, we will also sometimes refer to the transition function simply as the transition dynamics. For a given state and action, the transition is conditionally independent of all previous states and actions, which is known as the Markov property and the reason for the naming as an MDP. The value ${R}(s,a)\in\mathbb{R}$ provides an immediate evaluation after performing action $a$ in state $s$, which is also called the (instantaneous) reward. It is also possible that the instantaneous reward is 0 for some states, and one only receives a reward in specific states, for example, in so-called terminal states for which the transition function is zero. When both the state space $\mathcal{S}$} and the action space $\mathcal{A}$} are finite, we call the MDP a tabular MDP.
In an MDP, an $H$-step trajectory $\tau$ is a sequence of $H\in\mathbb{N}\setminus\{0\}$ state-action pairs ending in a terminal state. Formally, it is given by $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H})$. A trajectory $\tau$’s return ${R}(\tau)$ is the accumulated (discounted) rewards collected along this trajectory:
@@
Note that we here use the same notation for the return and the reward function; however, both have different signatures (trajectory vs. state-action pair). We can also define the return ${R}(\sigma)$ of a segment $\sigma$ in a similar manner. The return is well defined even if the horizon $H$ is infinite as long as $\gamma<1$. If the MDP is a tabular MDP and any trajectory has finite length, i.e., $H$ is necessarily finite, we call the MDP finite and otherwise infinite.
EnvironmentDynamicsObjectiveAgentAction $a_{t}$State $s_{t+1}$Reward $r_{t+1}$
@@
(b) RLHF with reward modeling.
Figure 1: Contrasting the standard RL setting with RLHF in its most common formulation, using a reward model. In each step, the policy commits to an action $a_{t}$ and receives the next state $s_{t+1}$ and either the true reward $r_{t+1}$ or an estimate $\hat{r}_{t+1}$ in return. In contrast to the standard RL setting, the true reward function is not known in the RLHF setting but instead learned from human feedback. This reward learning process is decoupled from policy learning and can happen fully asynchronously. The dataset consists of a set of queries $q_{i}$ (e.g., pairs of trajectory fragments) and their labels $l_{i}$ (e.g., a preference for one of the fragments).
@@
The basic interaction loop is depicted in Fig.˜1(a): The agent chooses an action $a_{t}\sim\pi(s_{t})$ based on its policy and the current state. As a consequence, the environment transitions to the new state $s_{t+1}\sim{P}(s_{t},a_{t})$, governed by the transition dynamics. The agent observes this new state and the reward $r_{t+1}=R(s_{t},a_{t})$, after which the interaction cycle repeats.
@@
where the expectation is with respect to policy $\pi$, transition function P{P}, and initial distribution $d_{0}$. To solve this problem, two families of RL approaches have been considered: model-based RL and model-free RL. The methods in the first family learn a model (i.e., P,R{P},{R}) of the underlying MDP to help solve the RL problem, while the methods in the second directly try to obtain a good policy without learning an MDP model. The second family can be further decomposed into two main categories: value-based methods and policy search methods. In deep RL, both value functions and policies are approximated with neural networks.
@@
where $s_{0}=s$ and $a_{0}=a$, and in the expectation, $a_{h}\sim\pi(\cdot\mid s_{h})$ as well as $s_{h}\sim{P}(\cdot\mid s_{h-1},a_{h-1})$ for $h\in[H-1]$. A policy can be naturally designed from a $Q$-function by choosing an action in a greedy manner in each state: $\pi(s)=\arg\max_{a}Q(s,a)$. Note that for a deterministic optimal policy $\pi^{*}$ it holds that $J(\pi^{*})=\mathbb{E}_{d_{0}}[Q^{*}(s,\pi^{*}(s))]$.
@@ 2.2 Preference-Based MDPs
In contrast to standard RL as described in the previous section, RLHF does not assume that a reward signal is available. It instead assumes the existence of an oracle (e.g., human labeler) that can provide information about the reward in a specific indirect manner. More precisely, in RLHF, the agent can make queries $q_{i}$ to the oracle, which in practice means asking for human feedback, and in response, the agent receives a label $l_{i}$, which in general gives a hint about the reward. In principle, the query can be made asynchronously to the actual conventional RL cycle. See Fig.˜1(b) for an illustration.
@@ 2.3 Reward Learning
RLHF approaches can be divided into two categories, depending on whether a utility-based approach is used for reward modeling or whether an alternative criterion that is detached from a utility concept is used (Gilbert et al., 2016; Gilbert & Weng, 2016; Wirth et al., 2017). Most works fall into the first category, on which this overview focuses. Such approaches assume a human-dependent utility function that can be used as a reward function in order to apply standard RL methods. Next, we will describe the commonly used approach for reward learning for the common setting of binary trajectory comparisons.
@@
In the context of RL, since ${R}_{\psi}(\tau)=\sum_{h=0}^{H-1}\gamma^{h}R_{\psi}(s_{h},a_{h})$, (2) can then directly be used to train a function approximator (e.g., single or ensemble of neural networks) to approximate $R$.
This approach accommodates the case of a noisy or unreliable oracle, in which case the Bradley-Terry model can be understood as the generative model of the oracle’s answers (or labels provided by the human labeler). When the oracle is reliable, more direct methods based on preference elicitation for recovering the reward function have been studied (Regan & Boutilier, 2009; 2011; Weng & Zanuttini, 2013; Gilbert et al., 2015; Sadigh et al., 2017; Wilde et al., 2018). In this survey, we will focus on the general case where the oracle may be noisy.
Note that, in contrast to the typical way of preference learning, the learned reward function is used to train an RL agent and not directly to compare trajectories. This discrepancy between the objective function used in reward learning and how the learned rewards are actually used may lead to suboptimal policies (Lindner et al., 2021b).
@@ 2.4 Reinforcement Learning from Human Feedback
In the RLHF setting as illustrated in Fig.˜1(b), the learning agent needs to solve an RL task without having access to a reward function. To this end, the agent usually simultaneously learns an approximation of the reward function (via the assumed utility function) and an RL policy. Therefore, a generic RLHF algorithm consists of repeating two phases: (1) reward learning and (2) RL training. The first phase can itself be decomposed into two main steps: (i) generate queries to ask the oracle, (ii) train a reward function approximator with the answers provided by the oracle. The RL training part is more conventional and is usually directly based on running a deep RL algorithm using the currently trained reward function approximator.
Algorithm 1 Generic RLHF Algorithm in an Actor-Critic Scheme.
@@
7:   Update $\psi$ using $\mathcal{D}$ (e.g., to maximize Eq.˜2)
@@
This basic generic algorithm is summarized in Algorithm˜1, where an off-policy actor-critic scheme is assumed to be used for the RL training part, but other RL policy learning approaches can, of course, also be used here. For an on-policy algorithm, such as PPO (Schulman et al., 2017), only the recently generated transitions are used for training. For a DQN-like algorithm, lines 9 to 11 would be replaced by a loop where transitions are generated by a behavior policy based on the current estimate of the $Q$-function (e.g., $\varepsilon$-greedy algorithm) and the $Q$ network is updated using mini-batches of transitions sampled from the replay buffer $\mathcal{B}$.
@@
  • The oracle may provide various types of feedback (see Section˜3). The questions of what information some given feedback provides and how observed feedback can be exploited need to be answered (see Section˜5).
  • Informative queries need to be generated to minimize the efforts of the oracle, which is crucial when it is a human (see Section˜4). Active learning techniques (see next subsection) can be adapted to face this challenge.
  • The RL agent is actually trained in a non-stationary environment since the reward approximator is concurrently updated. The RL training part therefore needs to account for this factor, e.g., using non-vanishing learning rates (see Section˜6).
  • There is also the question of how the agent’s performance can be meaningfully evaluated, especially if the reward function is not known (see Section˜8).
  • Collecting feedback directly from humans introduces its own challenges, such as the question of a suitable user interface and the associated issues of delay between query and feedback observation, or the feedback variability and reliability (see Section˜4). This may explain why many studies evaluate novel RLHF algorithms with simulated feedback.
A standard RL algorithm can be run in the RL training part, as done in most previous work in RLHF (although this may not be the best approach). This suggests that any improvements in a standard deep RL method (e.g., auxiliary losses (Gelada et al., 2019), planning in learned model (Hafner et al., 2020), curriculum learning (Narvekar et al., 2020), or data augmentation (Laskin et al., 2020; Lee et al., 2020; Lin et al., 2020b)) may potentially be transferred to the RLHF setting. In addition, most previous work in RLHF directly uses trajectories stored in replay buffer $\mathcal{B}$ to synthesize queries. An interesting research direction to explore in RLHF would be to specifically generate trajectories in order to be able to synthesize more informative queries (instead of only generating trajectories that are beneficial for RL training). This would lead to tackling a novel exploration-exploitation dilemma: Shall we visit state-action pairs that may be bad but may help better learn the reward function, or shall we visit state-action pairs that we currently think are good? This is further discussed in Section˜5.5.3.
@@ 2.5 Active Learning
In active learning (Settles, 2012), the task is to strategically select data points for labeling to minimize the amount of labeled data required to achieve a desired level of performance, which is particularly valuable in scenarios like RLHF where labeling is costly. Unlike batch learning, where labeled data is predetermined, active learning empowers the learner to actively select the most informative unlabeled instances for labeling, maximizing the learning process with limited labeled data. We will only briefly introduce the active learning task here and then discuss the strategies for creating informative queries considered thus far in Section˜4.
@@
At a high level, the main idea in active learning is to query data points to quickly reduce the epistemic (i.e., reducible) uncertainty about the predictions of the learned model, although other aspects can be important, such as the representativeness of the queried data points (see Wang et al. (2024b) for a survey). Two main representations are considered to describe this epistemic uncertainty: either using an ensemble of models or using a Bayesian representation. In both cases, a first basic approach selects queries using uncertainty-based criteria in order to focus on instances with high prediction uncertainty as measured by, e.g., variance or entropy computed over predictions. In contrast to the first approach, where the selection criteria are instance-based, a second approach considers criteria that may depend on all instances. Possible options are, for instance, expected model change, expected error reduction, or density-weighted uncertainty-based criteria. Here, the weights in the expectation or density allow us to take into account the distributional information about the instances and, therefore, to focus on the higher-density regions.
@@ 3 Feedback
Figure 2: RLHF diagram highlighting components discussed in this section.
Feedback mechanisms are fundamental to the success of any RL system. In the standard setting as described in Section˜2.1, RL agents expect feedback in the form of scalar immediate rewards. These rewards are most commonly determined by a hand-engineered reward function, which can be used to evaluate any state-action combination. As discussed in Section˜1.1, it is desirable to allow humans to refine behavior interactively through feedback instead of requiring them to pre-specify a reward function.
While a human could, in principle, assign rewards to each of the agent’s actions directly, thereby taking the role of the reward function, this is usually impractical for multiple reasons. The main challenge is the human effort required to provide rewards on a sufficiently regular basis, i.e., at least once per episode. In addition to that, directly integrating human rewards into the RL loop would require these rewards immediately, which would impede the learning pace while waiting for human feedback. Finally, the standard RL setting expects numeric state-action rewards, which are challenging to provide in a consistent manner.
In contrast to directly rewarding each of the agent’s actions, RLHF as discussed in this survey (see Section˜1.3) harnesses indirect and asynchronous feedback methods. Such methods avoid the challenges of immediate numeric rewards and are also better aligned with human interaction patterns, resulting in improved learning progress and human user experience.
@@
The remainder of this section will start by discussing relevant attributes of feedback types that can be used to classify them (Section˜3.1). We will then discuss common classes and examples of interactive feedback types (Section˜3.2) as well as some non-interactive types that can serve as initializations (Section˜3.3).